Two spherical objects are separated by a distance of 1.06 × 10-3 m. The objects are initially electrically neutral and are very small compared to the distance between them. Each object acquires the same negative charge due to the addition of electrons. As a result, each object experiences an electrostatic force that has a magnitude of 10.0 × 10-21 N. How many electrons did it take to produce the charge on one of the objects?
Answer #1

The electric force between the objects can be expressed as:

Fe = kq2/r2 where
k is electrostatic constant 9*109 N.m²/C²
q is the charge of each object
r is the distance separating them

So
q = (Fe.r2/k)0.5
q = (10*10-21*(1.06*10-3)2/9*109)0.5

=1.117*10-18 (with minus sign since it is negative)

One electron has charge
e = 1.6*10-19 C

So this charge represents q/e electrons
q/e = 1.117*10-18/1.6*10-19

= 7 electrons